Parth (parthkohli):

Try this sweet physics question. I'm not bragging that I was able to solve this, but I am.

Parth (parthkohli):

A certain mass of an ideal gas is enclosed in a cylinder of volume \(V_0\) fitted with a smooth heavy piston of mass \(m\) and cross-sectional area \(A\). The piston is displaced through a small distance downwards so as to compress the gas isothermally, and then it is let go. Take the atmospheric pressure as \(P\). Find the time period of the resultant oscillations.

OpenStudy (alekos):

I've never seen a physics question like this before!! Let me think about it

OpenStudy (kainui):

This sounds like a perpetual motion device lol

Parth (parthkohli):

That is because it is in ideal conditions. But if we extract energy, it should stop its motion, so it isn't a perpetual motion in that sense (you cannot extract infinite energy).

OpenStudy (irishboy123):

start with the ideal gas law \(pV = nRT = const \) as isothermal do \(d(pV) = 0 = pdV + V dp \implies dp = -\frac{p dV}{V}\) from there, for ***small*** displacement x, the restoring "linear" force for shm is \(dp A\), so \( |dp A| = - m \ddot x\). etc

OpenStudy (michele_laino):

I think that the system can be modeled with the subsequent 2-nd order ODE \[\Large M\frac{{{d^2}\left( {PS} \right)}}{{d{t^2}}} = - \left( {P - {P_0} - Mg} \right)S\] where \(S\) is the surface of the piston, \(M\) is its mass, and \(P_0\) is the atmospheric preesure

OpenStudy (michele_laino):

oops..pressure*

OpenStudy (alekos):

I get f''[x(t)] + xPA^2/mV = 0 where P = p + mg/A (p = atmospheric pressure)

OpenStudy (alekos):

looks similar to Michele's

OpenStudy (alekos):

and for the period i get T = 2πsqrt(mV/PA^2)

OpenStudy (alekos):

how does that look?

OpenStudy (michele_laino):

oops.. I have made a typo, here is the right equation  \[\Large \frac{{{d^2}z}}{{d{t^2}}} = \frac{{ - \left( {P - {P_0} - Mg} \right)S}}{M},\quad P = \frac{{nRT}}{{Sz}}\]

OpenStudy (anonymous):

All I have to say is that being able to solve a question like this is my goal. Good job!

OpenStudy (michele_laino):

more precisely, we can write this: \[\Large P = \frac{{nRT}}{{Sz}},\quad {P_0} + Mg = \frac{{nRT}}{{S{z_0}}}\] and for little oscillations, namely \[\huge z \approx {z_0}\] we have: \[\huge \ddot z + \frac{{nRT}}{{Mz_0^2}}\left( {z - {z_0}} \right) = 0\] therefore, the period of little oscillations is: \[\huge T = 2\pi \left( {\frac{{{V_0}}}{S}} \right)\sqrt {\frac{M}{{nRT}}} \]

OpenStudy (alekos):

so @parthkohli what's the answer

OpenStudy (alekos):

he's disappeared and left us holding the baby!

Parth (parthkohli):

Hey I'm back!

Parth (parthkohli):

Basically you're not given \(n, T\) but I'm sure you can use the equation of state.

Parth (parthkohli):

@alekos looks like you're right!

OpenStudy (alekos):

Awesome! thanks @parthkohli great question!

OpenStudy (michele_laino):

we can make this substitution: \[\huge nRT = {P_0}{V_0}\] so, we get: \[\huge T = 2\pi \sqrt {\frac{{M{V_0}}}{{{P_0}{S^2}}}} \]

Parth (parthkohli):

Substitute \(P = P_{atm} + mg/A\) for the actual final answer though.